OpenStudy (anonymous):

ive been working on this for so long, id really appreciate some help. An open-topped aquarium holds 40 cubic feet of water and is made of 60 square feet of glass. The length of the aquarium’s base is twice its width. Find the dimensions of the aquarium.

OpenStudy (anonymous):

Janes4, I would recommend that you start by setting up some equations.

OpenStudy (anonymous):

Such as \[V=40\].

OpenStudy (anonymous):

\[2lw+2lh+2wh=60\] (where l, w, and h are the lengths of the sides)

OpenStudy (anonymous):

and, \[l=2w\]

OpenStudy (anonymous):

so far i have 40=2(w^2)h and h=20/w^2

OpenStudy (mertsj):

\[lwh=40; l=2w; 2w(wh)=40; 2w^2h=40;w^2h=20\]

OpenStudy (joannablackwelder):

It looks like lw + 2lh + 2wh =60 to me since it is open-topped.

OpenStudy (mertsj):